Grade 12Algebra

Out of 1800 students in a school, 750 opted basketball, 800 opted cricket, and the remaining opted table tennis. If a student can opt only one game, find the ratio of
Number of students who opted for cricket to the number of students opting for basketball.

Pawan Prajapati
5 Years agoGrade 12

1 Answer

Harshit Singh
5 Years ago
Dear Student

Ratio of number of students who opted cricket to the number of students opting basketball
= 800 / 750 = 16 / 15

Thanks